OpenStudy (anonymous):

solve the proportion x/x+1=4/x+4

OpenStudy (mertsj):

\[\frac{x}{x+1}=\frac{4}{x+4}\]

OpenStudy (mertsj):

Cross Multiply

OpenStudy (mertsj):

\[x^2+4x=4x+4\]

OpenStudy (mertsj):

\[x^2=4\]

OpenStudy (mertsj):

\[x=\pm2\]

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

@ Mertsj How did you type the division(one above another)?

OpenStudy (mertsj):

Go to the equation below this box. Type: \frac{numerator}{denominator}

OpenStudy (anonymous):

Thanks

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

Hw do u add fraction here?